OpenStudy (anonymous):

6i(4 - i^18)(3 - 7i^63) heeeelllllppp

OpenStudy (anonymous):

What are you having trouble with?

OpenStudy (anonymous):

i think i got the answer but im not sure if its right. i got 72i - 168i^64 - 18i^19 + 42i^82

OpenStudy (anonymous):

Well, we can simplify that, right? what is i^64?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Right, how did you get that?

OpenStudy (anonymous):

Anyways, simplify your i terms first, then multiply out your expression.

OpenStudy (anonymous):

i divided 64 by 4 and used the little i circle thing to count around 16 times

OpenStudy (anonymous):

can i do that last or should do that first?

OpenStudy (anonymous):

It will make it easier, and you will have less chance of making a mistake if you simplify your i terms first.

OpenStudy (anonymous):

so i got a final answer of 72i-168 +18i -42

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

90i - 210 actually

OpenStudy (anonymous):

or -90 - 210 because when combining two i's it becomes i^2

OpenStudy (anonymous):

6i(4 - i^18)(3 - 7i^63)=6i(4-(-1))(3-(-i))=6i(5)(3+i)=30i(3+i)=90i-30

OpenStudy (anonymous):

So, not quite. You probably made a little arithmetic error somewhere.

OpenStudy (anonymous):

wait so how did you go from 3 - 7i^63 to 3 + i? what happened to the 7?

OpenStudy (anonymous):

I'm terribly sorry. I appear to have made the error.

OpenStudy (anonymous):

dfyodor, did a boo boo.

OpenStudy (anonymous):

90i-210*

OpenStudy (anonymous):

Ammend the above to read: 6i(4 - i^18)(3 - 7i^63)=6i(4-(-1))(3-7(-i))=6i(5)(3+7i)=30i(3+7i)=90i-210

OpenStudy (anonymous):

okay awesome, got it!